# GC-xLSTM: Neural Granger Causality with xLSTMs

### v1

@@ 3.2 Optimizing for Strict Input Sparsity
23:$\mathcal{E}\leftarrow\left\{(v,w)\in\mathcal{V}\times\mathcal{V}\;\middle|\;\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}>0\right\}$
24:return extracted graph $(\mathcal{V},\mathcal{E})$
Figure 2: The optimization procedure for GC-xLSTM. GC-xLSTM alternates between GD and proximal GD.
The purpose of the feature selector $\bm{W}$ is to only “pay attention” to as many variates as necessary for successful forecasting. A common approach to achieve this sparsity on the Granger causal relationships is via the lasso regularization explained in Section 2.2. We use the group lasso penalty [Yuan and Lin, 2006, Simon and Tibshirani, 2012] on the initial projection layer of GC-xLSTM as a structured sparsity-inducing penalty that encourages the selection of entire groups of variables, encoded as the columns of $\bm{W}_{v}$. We adaptively compress by learning a reduction coefficient $\bm{\alpha}_{v}\in\mathbb{R}^{V}$ that selects which of the $V$ columns of $\bm{W}_{v}$ are redundant. We perform this compression of $\bm{W}_{v}$ in a joint procedure with the general optimization of the forecasting model, as provided in Algorithm 1. Specifically, we perform two updates per optimization step for each of the variates, as Figure 2 depicts. Firstly, we optimize the projection weights $\bm{\phi}$, the reduction coefficients $\bm{\alpha}$, and the xLSTM parameters $\bm{\theta}$ using mini-batch gradient descent. This corresponds to 12 to 18 in Algorithm 1. This optimizes the following loss expected over the time series data $\bm{S}$:
Note that we, crucially, only descent on the reduction coefficient $\bm{\alpha}$ in $\mathcal{L}_{\text{red}}$ an not on $\bm{W}_{v}$, as the stop-gradient $\operatorname{sg}(\cdot{})$ denotes. This sparsity optimization is instead performed by the second step in the procedure, shown in 19, where a proximal gradient descent step is taken to dynamically shrink $\bm{W}_{v}$ proportional to $\bm{\alpha}_{v}$. The $(\cdot)_{+}$ denotes truncation as $\max(\cdot,0)$. The compression update takes a descent step towards the gradient of
followed by a soft thresholding. Intuitively, the $\mathcal{L}_{\text{red}}$ component of Eq. 12 keeps $\bm{W}_{v}$ fixed while learning $\bm{\alpha}_{v}$, and Eq. 13 keeps $\bm{\alpha}_{v}$ fixed in the proximal step to compress $\bm{W}_{v}$. Figure 3 depicts the intuition of the proximal gradient step and soft-thresholding in 19.
Details on learning the reduction loss $\mathcal{L}_{\text{red}}$.
It is worth briefly discussing the use of the logarithm in Eq. 12. It mainly gives more equal weight to the decreases in $\bm{W}_{v}$ column norms and encourages learning of better sparse Granger causal relations. It furthermore normalizes the gradient updates to $\bm{\alpha}_{v}$. Empirically, this loss engineering allowed training models that were significantly more robust to noise and changes to the sparsity hyperparameter $\lambda$. This was reflected by a more stable variable usage and predictive loss $\mathcal{L}_{\text{pred}}$.
Ensuring non-negativity of the reduction coefficients $\bm{\alpha}$.
For the proximal update step to be well-behaved, we need to ensure that $\bm{\alpha}_{v}$ results in a convex combination of column weights, i.e., that $\bm{\alpha}_{v}^{w}>0$ for all $w\in\mathcal{V}$ and $\bm{\alpha}_{v}^{T}\mathds{1}=1$. We achieve this by re-parameterizing it as $\bm{\alpha}_{v}=\mathrm{softmax}\left(\bm{\beta}_{v}\right)$, and learning $\bm{\beta}_{v}$ instead of $\bm{\alpha}_{v}$.
Figure 3: The intuition of the compression step in 19. Lighter colors indicate matrix columns with smaller norms. In some cases, the matrix retains the high values, indicating its importance in uncovering the Granger causal relations while thinning out others.
Intuitive dynamics of the gradient update step.
The weights $\bm{W}_{v}$ in the reduction loss $\mathcal{L}_{\text{red}}$ of Eq. 12 only serve to learn good reduction coefficients $\bm{\alpha}_{v}$. Deriving the gradient of the penalty term $\mathcal{L}_{\text{reg}}$ with respect to the underlying $\bm{\beta}$ provides a helpful intuition of the training dynamics:
$\frac{\partial}{\partial\beta_{v}^{w}}\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}$ $=\frac{\partial}{\partial{\beta_{v}^{w}}}\sum_{w=1}^{V}\mathrm{softmax}(\bm{\beta}_{v})^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}$
$=\alpha_{v}^{w}\left(\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}-\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}\right)$
$\Longrightarrow\frac{\partial}{\partial\beta_{v}^{w}}\mathcal{L}_{\text{reg}}$ $=\lambda\frac{\partial}{\partial\beta_{v}^{w}}\log\left(\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}\right)$
$=\lambda\alpha_{v}^{w}\frac{\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}-\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}}{\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}}$
$=\lambda\alpha_{v}^{w}\left(\frac{\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}}{\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}}-1\right)$
Table 1: Comparison of AUROC for Granger causality selection among different approaches, as a function of the forcing constant $F$ and the length of the time series $T$. Bold denotes best, underlined denotes the runner-up.
Model $F=10$ $F=40$
$T=250$ $T=500$ $T=1000$ $T=250$ $T=500$ $T=1000$
cMLP 86.6 ± 0.2 96.6 ± 0.2 98.4 ± 0.1 84.0 ± 0.5 89.6 ± 0.2 95.5 ± 0.3
cLSTM 81.3 ± 0.9 93.4 ± 0.7 96.0 ± 0.1 75.1 ± 0.9 87.8 ± 0.4 94.4 ± 0.5
GC-KAN 86.9 ± 0.5 92.1 ± 0.3 96.2 ± 0.2 86.3 ± 0.2 87.1 ± 0.4 95.7 ± 0.2
GC-xLSTM (ours) 99.6 ± 0.2 99.3 ± 0.3 99.1 ± 0.2 71.1 ± 0.3 88.0 ± 0.2 91.4 ± 0.2
Practical Considerations.
Furthermore, we perform staged optimization of $\bm{\alpha}$, which is initialized to a uniform distribution by setting all $\bm{\beta}=0$. We only start training $\bm{\alpha}$ after exploring the prediction loss and having obtained a reasonably compressed forecaster, which is controlled by the hyperparameter $K$ in Algorithm 1 (see 16). While we present the method with mini-batch gradient descent for conciseness, improved convergence can be achieved with more modern optimizers, such as Adam [Kingma and Ba, 2017].
4 Experimental Evaluation
We conduct extensive experiments on three datasets to assess the practical effectiveness of GC-xLSTM. We will now explain the underlying parameters used to train the GC-xLSTM architecture and then discuss the used data sets before presenting our obtained results in detail.
4.1 Model Details
For our component-wise networks, we use a single xLSTM block comprised of one sLSTM layer, followed by a linear layer to predict the next step of the modeled variate. The hidden dimension of the sLSTM layer is set to 32 for all datasets. We find that the presence of a gated MLP to up- and down-project the hidden states of the sLSTM layer does not significantly improve the performance of the model, so we choose to omit it in all our experiments for simplicity. We do not use any mLSTM layers in our experiments, as we find that the sLSTM layer is sufficient to capture all relevant long-range dependencies in the data.
We use the Adam optimizer with a weight decay of 0.1 for training. We schedule the learning rate to follow a linear warmup of 2,000 iterations to $10^{-4}$, followed by cosine annealing until the end of training for a total of 13,000 steps. We start learning the reduction coefficients after a warmup period of $K=1,500$ iterations, during which the uniform compression across all columns combined with the prediction loss gives reasonable priors for the gradient directions of the reduction coefficients. Due to the moderate size of the datasets, we performed full-batch gradient descent. Only the sparsity hyperparameter $\lambda$ was tuned specifically for each of the settings. We will describe them next.
4.2 Datasets
We now introduce the three datasets used to test the efficacy of GC-xLSTM.
@@ 4.2 Datasets
The $V$-dimensional Lorenz-96 model [Karimi and Paul, 2010] is a chaotic multivariate dynamical system governed by the differential equations
$\frac{dx_{t,i}}{dt}=(x_{t,i+1}-x_{t,i-2})x_{t,i-1}-x_{t,i}+F,$ (14)
where $F$ is the forcing constant that determines the degree of nonlinearity in the system. Low values of $F$ correspond to near-linear dynamics, while higher values induce chaotic behavior. Following the setting of Tank et al. [2021] we best comparability, we simulate $V=20$ variates with a sampling rate of $\Delta t=0.05$ for a total of $T\in\{250,500,1000\}$ time steps. We use two forcing constants $F\in\{10,40\}$ to test our model under different levels of non-linearity.
@@
The Moléne dataset [Girault, 2015] contains hourly temperatures recorded by sensors at $V=32$ locations in Brittany, France during $T=744$ hours. The objective is to understand the spatio-temporal dynamics of the temperature and to assess the extent to which the model can uncover complex relationships in weather by considering only local observations and their geographical positioning.
(a) $\lambda$ = 8.
(b) $\lambda=10$.
Figure 4: Granger causal relations on the Moléne Dataset. GC-xLSTM uncovers several intersting Granger causal relations giving us an insight into the dynamic weather patterns. We can observe that the sparsity of the learned Granger causal relations increases with the value of $\lambda$.
@@
We next apply our methodology to detect complex, nonlinear dependencies in human motion capture (MoCap) recordings. In contrast to the Lorenz-96 simulated dataset results, this analysis allows us to more easily visualize and interpret the learned network. We consider a data set from the CMU MoCap database [CMU, 2009]. The data recorded from 30 unique regions in the body consists of $V=54$ joint angle and body position recordings across multiple different subjects. Since some regions, like the neck, have multiple degrees of freedom in both translation and rotation, we consider the GC relations between two joints based on edges between all movement directions. The motion ranges from locomotion (e.g., running and walking) over physical activities such as gymnastics and dancing to day-to-day social interactions.
(a) Salsa with $\lambda=4$.
(b) Salsa with $\lambda=5$.
(c) Running with $\lambda=4$.
(d) Running with $\lambda=6$.
Figure 5: GC-xLSTM captures complex human motions. GC-xLSTM is able to uncover complex real-world dependencies in the Human Motion Capture dataset giving us an intiuative understanding of the learned interactions.
4.3 Experimental Results
Lorenz-96.
We compare the results of GC-xLSTM with three baseline methods: (1) the cMLP model, (2) the cLSTM model, both proposed by Tank et al. [2021], and (3) GC-KAN, proposed by Lin et al. [2024]. The baseline results are taken as reported in the original papers. Table 1 shows the AUROC scores for the Granger causality selection task on the Lorenz-96 dataset. GC-xLSTM significantly outperforms all the baseline methods for the $F=10$ case, across various time series. This suggests that GC-xLSTM is able to capture the underlying Granger causal relationships in the presence of limited data more efficiently than other methods. The more focused compression of norms of columns in the projection layer combined with the superior modeling power of xLSTM enables the GC-xLSTM to selectively squash false positives while retaining true positives. For $F=40$, GC-xLSTM performs competitively to the baselines in the presence of high non-linearity but is outperformed for $T=250$. We postulate this is because the learning of reduction coefficients does not have access to high-quality projection weight priors during the initial stages of training, due to less data. GC-xLSTM substantially overcomes this challenge and displays similar improvement trends as the baselines with increasing time steps. Furthermore, the model displays strong prediction accuracies as seen in Figure 6.
Moléne.
Unlike Zoroddu et al. [2024], who incorporate graph prior knowledge based on sensor locations, our approach learns the GC structure solely from the temperature observations. This ensures that GC-xLSTM does not inherently favor regional connections over long-range dependencies, allowing it to discover dominant weather patterns operating both locally and across broader spatial scales. Figure 4 shows the obtained Granger causal graphs under different sparsity constraints. The dense structure of 4(a) exhibits a richer set of Granger causal interactions, while the sparser graph in 4(b) highlights the most significant edges. By adjusting $\lambda$, we can balance granularity and interpretability, allowing for insights into both local and large-scale temperature dependencies.
Human Motion Capture.
To analyze GC relations between body joints, we focus on two specific activities: Salsa dancing and running, which provide interpretable motion patterns. Figure 5 shows the results of the learned graphs for those activities. A closer look at the subfigures offers us an intuitive understanding of the learned interactions. For example, in the Salsa dance, we observe edges from the feet to the knees and the arms, supporting the characteristic movements of the lower driving the upper body. We can also see the cross-limb correlation, with movements initiating on one side of the body and being propagated to the other. Similarly, the results for running strongly establish the lower limbs as primary motion drivers, with edges from the feet and knees to the arms. We also see cyclic dependencies between the knees, ankles, and feet, capturing the repetitive, alternating nature of the gait.
4.4 Inspecting Training Dynamics
Here, we elaborate on the utility of the logarithm in the reduction loss $\mathcal{L}_{\text{red}}$ of Eq. 12. The logarithm term incentivizes the model to explore sparser solutions to Granger causal discovery by allowing the training to move forward over any local minima that make use of the full set of input variates. As the projection matrix becomes sparser and the input variates vanish from consideration by the model, the prediction loss $\mathcal{L}_{\text{pred}}$ will slightly increase. As seen in Figure 7, an increase in the sparsity of the feature selectors $\bm{W}$ drives down the loss and enables the model to learn more meaningful Granger causal relations.

### v2

@@ Appendix A Full Algorithm
21:$\mathcal{E}\leftarrow\left\{(v,w)\in\mathcal{V}\times\mathcal{V}\;\middle|\;\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}>0\right\}$
22:return extracted graph $(\mathcal{V},\mathcal{E})$
Appendix B On the Approximation Capabilities of sLSTM blocks
Classic RNNs have long been known to be extremely powerful models of computation. Specifically, they are Turing complete [Siegelmann and Sontag, 1995] and, by extension, universal function approximators [Song et al., 2023]. Traditional LSTMs as proposed by Hochreiter and Schmidhuber [1997] are universal function approximators being at least as powerful as the RNNs [Song et al., 2023, Corollary 16]. Due to their novelty, the sLSTM cells as presented in Section 2.3 have not yet received the same degree of theoretical analysis. Yet, it appears natural to extend the same reduction to RNNs as has been shown for LSTMs, since the main technicality is the differing normalization of the hidden state [Beck et al., 2024, Eq. 36] and the alternative exponential activation function, which Song et al. [2023] abstract away with. sLSTM blocks then wrap these cells with additional operations (cf. Beck et al. [2024], App. A.4). They, however, can be carefully configured to reduce to only rescaling Layer Normalizations and residual connections, which are known to, again, yield universal function approximators for many architectures [Tabuada and Gharesifard, 2020, Marchi et al., 2021]. This can be shown by first omitting the anyway optional upfront convolution and Swish activation. We, furthermore, set the number of heads to $N_{h}=1$, causing the per-head Group Normalization after the cell to degenerate to yet another Layer Normalization. Lastly, the final post up-projection of $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$ is defined as
$\bm{x}_{\text{out}}=\bm{W}_{3}\left(\left(\bm{W}_{1}\bm{x}_{\text{in}}+\bm{b}_{1}\right)\odot\operatorname{GeLU}\left(\bm{W}_{2}\bm{x}_{\text{in}}+\bm{b}_{2}\right)\right)+\bm{b}_{3}.$
With the right instantiation, where $\alpha$ is chosen such that $\operatorname{GeLU}(\alpha)=1$, we obtain
$\bm{x}_{\text{out}}$ $=\begin{pmatrix}\mathbb{I}&0&0&0\\ 0&\mathbb{I}&0&0\\ 0&0&\mathbb{I}&0\end{pmatrix}\left(\left(\begin{pmatrix}\mathbb{I}&0&0\\ 0&\mathbb{I}&0\\ 0&0&\mathbb{I}\\ 0&0&0\end{pmatrix}\bm{x}_{\text{in}}+\bm{0}\right)\odot\operatorname{GeLU}\left(\bm{0}\bm{x}_{\text{in}}+\alpha\bm{1}\right)\right)+\bm{0}$
$=\begin{pmatrix}\mathbb{I}&0&0&0\\ 0&\mathbb{I}&0&0\\ 0&0&\mathbb{I}&0\end{pmatrix}\begin{pmatrix}\mathbb{I}&0&0\\ 0&\mathbb{I}&0\\ 0&0&\mathbb{I}\\ 0&0&0\end{pmatrix}\bm{x}_{\text{in}}\odot\bm{1}$
$=\bm{x}_{\text{in}}.$
This, again, shows that sLSTM blocks are at least as general as LSTMs. While a rigorous proof is far beyond the scope of this work, this investigation still underpins the strong capabilities of this architecture. These theoretical considerations align well with empirical findings, showing that xLSTM blocks are at least as effective as LSTMs [Beck et al., 2024].
Appendix C Dataset Details
This section details all six datasets used in our empirical evaluation of GC-xLSTM.
@@ Appendix C Dataset Details
The $V$-dimensional Lorenz-96 model [Karimi and Paul, 2010] is a chaotic multivariate dynamical system governed by the differential equations $\frac{dx_{t,i}}{dt}=(x_{t,i+1}-x_{t,i-2})x_{t,i-1}-x_{t,i}+F,$ with the external forcing coefficient $F$ regulating the nonlinearity of the system. Low values of $F$ correspond to near-linear dynamics, while higher values induce chaotic behavior. There are two sources of randomness in the system. Firstly, we sample i.i.d. starting conditions from $\mathcal{N}(0,0.01)$. Secondly, in each step of the simulation, we add i.i.d. noise sampled from $\mathcal{N}(0,0.1)$ to $x_{t,i}$. Following the setting of Tank et al. [2022] for best comparability, we simulate $V=20$ variates with a sampling rate of $\Delta t=0.05$ for a total of $T=500$ time steps. We use two forcing constants $F\in\{10,40\}$ to test our model under different levels of non-linearity.
fMRI.
Discovering connectivity networks within (human) brains is a key application of GC detection methods. To this end, brain activity is measured non-invasively using functional magnetic resonance imaging (fMRI) over time and grouped into regions between which connections are to be found. Specifically, we used the realistic simulations of blood-oxygen-level-dependent (BOLD) deconvolved data of Smith et al. [2011]. The data contains $V=15$ variates over $T=200$ time steps.
@@
The Moléne dataset [Girault, 2015] contains hourly temperatures recorded by sensors at $V=32$ locations in Brittany, France, during $T=744$ hours. The objective is to understand the spatio-temporal dynamics of the temperature and to assess the extent to which the model can uncover complex relationships in weather by considering only local observations and their geographical positioning.
@@
We also apply our methodology to detect complex, nonlinear dependencies in human motion capture (MoCap) recordings. In contrast to the Lorenz-96 simulated dataset results, this analysis allows us to more easily visualize and interpret the learned network. We consider a data set from the CMU MoCap database [CMU, 2009]. The data recorded from 30 unique regions in the body consists of $V=54$ joint angle and body position recordings across multiple different subjects. Since some regions, like the neck, have multiple degrees of freedom in both translation and rotation, we consider the GC relations between two joints based on edges between all movement directions. The motion ranges from locomotion (e.g., running and walking) over physical activities such as gymnastics and dancing to day-to-day social interactions.
Company Fundamentals.
Another common field of application of GC detection methods is discovering links in economic data, where it is otherwise hard to maintain an overview. We, therefore, benchmark GC-xLSTM on company fundamentals data [Divo et al., 2024]. The dataset contains $V=20$ economic variables such as the Net Income and the Total Liabilities of 2527 companies. The data was collected quarterly from 2009 Q1 to 2023 Q3, resulting in only $T=56$ time steps.
VAR.
Following the well-known setup of Tank et al. [2022], we generate a two-step linear auto-regressive process in $V\in\{10,20\}$ variates. All variables depend on themselves, and an additional three random dependencies are added as targets to be discovered. The number of time steps varies in $T\in\{250,500,1000\}$.
Appendix D Extended Experimental Results
This section supplements the experimental findings of Section 4.1. Specifically, Figure 7 provides the full set of relations extracted from the Company Fundamentals dataset.
Figure 7: GC-xLSTM uncovers relationships between economic variables of individual companies.
Appendix E Additional Insights into Training
Figure 8(a) shows how the different loss components change during training. The robustness of training is reflected in the stability of variable usage once an optimal set is found, as Figure 8(b) shows.
(a) Visualization of the different loss components.
(b) Variable drops once the threshold $k\geq K$ is reached and then quickly stabilizes.